# The Illusion of Progress: Re-evaluating Hallucination Detection in LLMs

### v1

@@ 3.6 Implementation Details
We utilize pretrained model weights from the Hugging Face Transformers (Wolf et al., 2020) without any additional fine-tuning. Following (Farquhar et al., 2024), we generate 10 samples ($n=10$) using temperature 1.01.01.0 for uncertainty estimation. Additionally, we generate one "best answer" sample with temperature 0.10.10.1 to serve as the best-generation estimate for performance evaluation.
@@ 5.1 Limitations of ROUGE for Factual Accuracy Assessment in QA
Model Metric NQ-Open SQuAD Trivia-QA Mean ROUGE LLM $\Delta$ROUGE LLM $\Delta$ROUGE LLM $\Delta$ROUGE LLM $\Delta$Llama Perplexity 0.709 0.700 -1.2 0.703 0.687 -2.4 0.733 0.789 7.2 0.715 0.725 1.2 LN-Entropy 0.521 0.605 13.9 0.558 0.611 8.7 0.563 0.636 11.5 0.547 0.617 11.4 SE 0.778 0.742 -4.8 0.707 0.705 -0.2 0.769 0.832 7.6 0.751 0.760 0.9 Eigenscore 0.816 0.686 -19.0 0.720 0.638 -12.7 0.752 0.734 -2.5 0.763 0.686 -11.4 eRank 0.825 0.632 -30.6 0.754 0.621 -21.4 0.717 0.660 -8.6 0.765 0.638 -20.2 LogDet 0.511 0.515 0.7 0.521 0.536 2.7 0.604 0.509 -18.6 0.545 0.520 -5.1 Mistral Perplexity 0.852 0.584 -45.9 0.516 0.500 -3.2 0.843 0.627 -34.4 0.737 0.570 -27.8 LN-Entropy 0.718 0.645 -11.3 0.734 0.657 -11.7 0.586 0.596 1.8 0.679 0.633 -7.1 SE 0.836 0.729 -14.7 0.784 0.701 -11.9 0.726 0.707 -2.6 0.782 0.712 -9.7 Eigenscore 0.873 0.669 -30.4 0.803 0.648 -24.0 0.775 0.652 -18.9 0.817 0.656 -24.4 eRank 0.925 0.678 -36.4 0.518 0.511 -1.3 0.851 0.645 -31.9 0.765 0.611 -23.2 LogDet 0.628 0.508 -23.6 0.562 0.518 -8.5 0.843 0.606 -39.2 0.678 0.544 -23.8
@@ 5.3 Impact of Few-Shot Examples on Evaluation Reliability
[…]
@@
[…]
@@ G.1 Length-Based Penalties
ROUGE systematically penalizes factually correct but verbose answers. In this example, despite providing accurate information with helpful context, the response receives a low score purely due to length mismatch. As shown in Figure 5, this bias affects longer responses regardless of their factual accuracy, with responses exceeding 100100100 tokens consistently scoring below our 0.30.30.3 threshold. Notably, this is the most frequent type of error ROUGE makes.
@@ H.2 Metric Evaluation: AUROC
Model Metric NQ-Open SQuAD Trivia-QA Mean ROUGE LLM $\Delta$ROUGE LLM $\Delta$ROUGE LLM $\Delta$ROUGE LLM $\Delta$Llama Perplexity 0.709 0.700 -1.2 0.703 0.687 -2.4 0.733 0.789 7.2 0.715 0.725 1.2 Llama LN-Entropy 0.521 0.605 13.9 0.558 0.611 8.7 0.563 0.636 11.5 0.547 0.617 11.4 Llama SE 0.778 0.742 -4.8 0.707 0.705 -0.2 0.769 0.832 7.6 0.751 0.760 0.9 Llama Eigenscore 0.816 0.686 -19.0 0.720 0.638 -12.7 0.752 0.734 -2.5 0.763 0.686 -11.4 Llama eRank 0.825 0.632 -30.6 0.754 0.621 -21.4 0.717 0.660 -8.6 0.765 0.638 -20.2 Llama Len 0.834 0.616 -35.3 0.777 0.622 -24.9 0.760 0.691 -10.0 0.790 0.643 -23.4 Llama LogDet 0.511 0.515 0.7 0.521 0.536 2.7 0.604 0.509 -18.6 0.545 0.520 -5.1 Llama Mean-Len 0.825 0.654 -26.1 0.743 0.643 -15.7 0.771 0.743 -3.8 0.780 0.680 -15.2 Llama Std-Len 0.711 0.644 -10.5 0.664 0.627 -6.0 0.759 0.754 -0.7 0.711 0.675 -5.7 Mistral Perplexity 0.852 0.584 -45.9 0.516 0.500 -3.2 0.843 0.627 -34.4 0.737 0.570 -27.8 Mistral LN-Entropy 0.718 0.645 -11.3 0.734 0.657 -11.7 0.586 0.596 1.8 0.679 0.633 -7.1 Mistral SE 0.836 0.729 -14.7 0.784 0.701 -11.9 0.726 0.707 -2.6 0.782 0.712 -9.7 Mistral Eigenscore 0.873 0.669 -30.4 0.803 0.648 -24.0 0.775 0.652 -18.9 0.817 0.656 -24.4 Mistral eRank 0.925 0.678 -36.4 0.518 0.511 -1.3 0.851 0.645 -31.9 0.765 0.611 -23.2 Mistral Len 0.934 0.634 -47.2 0.860 0.624 -37.8 0.929 0.673 -37.9 0.908 0.644 -41.0 Mistral LogDet 0.628 0.508 -23.6 0.562 0.518 -8.5 0.843 0.606 -39.2 0.678 0.544 -23.8 Mistral Mean-Len 0.890 0.643 -38.4 0.828 0.626 -32.2 0.875 0.667 -31.3 0.864 0.645 -34.0 Mistral Std-Len 0.516 0.512 -0.7 0.540 0.505 -6.9 0.613 0.572 -7.2 0.556 0.530 -4.9
@@
Model Metric NQ-Open SQuAD Trivia-QA Mean ROUGE LLM $\Delta$ROUGE LLM $\Delta$ROUGE LLM $\Delta$ROUGE LLM $\Delta$Llama Perplexity 0.814 0.767 -6.1 0.736 0.758 2.9 0.800 0.826 3.1 0.783 0.784 -0.0 Llama LN-Entropy 0.753 0.732 -2.9 0.663 0.717 7.5 0.799 0.829 3.6 0.738 0.759 2.7 Llama SE 0.738 0.730 -1.1 0.688 0.741 7.1 0.800 0.849 5.7 0.742 0.773 3.9 Llama Eigenscore 0.813 0.744 -9.3 0.725 0.733 1.2 0.745 0.762 2.3 0.761 0.746 -1.9 Llama eRank 0.794 0.714 -11.2 0.708 0.681 -4.0 0.620 0.638 2.8 0.707 0.678 -4.1 Llama Len 0.761 0.686 -10.9 0.694 0.687 -1.0 0.620 0.640 3.1 0.692 0.671 -2.9 Llama LogDet 0.729 0.690 -5.6 0.659 0.636 -3.7 0.590 0.618 4.5 0.659 0.648 -1.6 Llama Mean-Len 0.799 0.730 -9.4 0.713 0.716 0.4 0.681 0.716 4.8 0.731 0.721 -1.4 Llama Std-Len 0.777 0.727 -7.0 0.705 0.721 2.2 0.783 0.806 2.9 0.755 0.751 -0.6 Mistral Perplexity 0.804 0.632 -27.1 0.782 0.636 -23.0 0.744 0.637 -16.7 0.777 0.635 -22.3 Mistral LN-Entropy 0.727 0.619 -17.4 0.785 0.667 -17.7 0.750 0.692 -8.3 0.754 0.659 -14.5 Mistral SE 0.772 0.734 -5.3 0.737 0.698 -5.6 0.741 0.765 3.1 0.750 0.732 -2.6 Mistral Eigenscore 0.789 0.686 -15.0 0.775 0.691 -12.2 0.717 0.706 -1.5 0.760 0.694 -9.6 Mistral eRank 0.874 0.698 -25.1 0.829 0.690 -20.1 0.786 0.703 -11.8 0.830 0.697 -19.0 Mistral Len 0.879 0.664 -32.2 0.857 0.685 -25.1 0.858 0.729 -17.7 0.865 0.693 -25.0 Mistral LogDet 0.737 0.663 -11.2 0.687 0.631 -8.9 0.612 0.630 2.9 0.679 0.641 -5.7 Mistral Mean-Len 0.834 0.683 -22.1 0.822 0.705 -16.5 0.806 0.750 -7.4 0.821 0.713 -15.3 Mistral Std-Len 0.609 0.577 -5.6 0.629 0.589 -6.8 0.663 0.665 0.3 0.634 0.610 -4.0
@@ H.3 Metric Evaluation: PR-AUC
Model Metric NQ-Open SQuAD Trivia-QA Mean ROUGE LLM $\Delta$ROUGE LLM $\Delta$ROUGE LLM $\Delta$ROUGE LLM $\Delta$Llama Perplexity 0.833 0.680 -22.4 0.863 0.823 -4.8 0.594 0.514 -15.6 0.763 0.672 -14.3 Llama LN-Entropy 0.717 0.611 -17.4 0.793 0.773 -2.6 0.570 0.652 12.5 0.693 0.679 -2.5 Llama SE 0.845 0.695 -21.5 0.864 0.829 -4.1 0.575 0.533 -7.9 0.761 0.686 -11.2 Llama Eigenscore 0.850 0.670 -26.8 0.866 0.809 -7.1 0.565 0.574 1.6 0.760 0.684 -10.8 Llama eRank 0.782 0.607 -28.9 0.820 0.783 -4.6 0.674 0.760 11.3 0.759 0.717 -7.4 Llama Len 0.865 0.681 -27.2 0.885 0.820 -8.0 0.605 0.548 -10.4 0.785 0.683 -15.2 Llama LogDet 0.852 0.659 -29.2 0.873 0.810 -7.8 0.602 0.562 -7.1 0.776 0.677 -14.7 Llama Mean-Len 0.851 0.658 -29.3 0.870 0.808 -7.7 0.573 0.568 -0.9 0.765 0.678 -12.6 Llama Std-Len 0.825 0.647 -27.6 0.846 0.802 -5.5 0.562 0.570 1.4 0.744 0.673 -10.6 Mistral Perplexity 0.664 0.536 -23.8 0.951 0.754 -26.0 0.690 0.752 8.3 0.768 0.681 -13.8 Mistral LN-Entropy 0.882 0.664 -32.8 0.920 0.790 -16.4 0.625 0.633 1.3 0.809 0.696 -16.0 Mistral SE 0.956 0.725 -31.8 0.964 0.819 -17.7 0.808 0.510 -58.3 0.909 0.685 -35.9 Mistral Eigenscore 0.957 0.698 -37.1 0.965 0.804 -20.0 0.818 0.544 -50.3 0.913 0.682 -35.8 Mistral eRank 0.658 0.506 -30.0 0.955 0.755 -26.4 0.534 0.704 24.2 0.716 0.655 -10.7 Mistral Len 0.964 0.682 -41.4 0.973 0.803 -21.0 0.849 0.536 -58.4 0.929 0.674 -40.3 Mistral LogDet 0.964 0.699 -37.9 0.950 0.753 -26.1 0.847 0.550 -54.1 0.920 0.667 -39.4 Mistral Mean-Len 0.958 0.671 -42.8 0.966 0.786 -22.9 0.833 0.548 -52.1 0.919 0.668 -39.3 Mistral Std-Len 0.891 0.583 -52.8 0.889 0.724 -22.7 0.755 0.605 -24.8 0.845 0.637 -33.4
@@
Model Metric NQ-Open SQuAD Trivia-QA Mean ROUGE LLM $\Delta$ROUGE LLM $\Delta$ROUGE LLM $\Delta$ROUGE LLM $\Delta$Llama Perplexity 0.844 0.824 -2.4 0.861 0.891 3.4 0.551 0.502 -9.8 0.752 0.739 -2.9 Llama LN-Entropy 0.810 0.796 -1.8 0.828 0.874 5.3 0.525 0.522 -0.5 0.721 0.731 1.0 Llama SE 0.814 0.802 -1.5 0.842 0.879 4.3 0.536 0.506 -6.1 0.731 0.729 -1.1 Llama Eigenscore 0.829 0.802 -3.4 0.852 0.876 2.7 0.511 0.542 5.7 0.731 0.740 1.7 Llama eRank 0.746 0.726 -2.8 0.711 0.762 6.8 0.679 0.737 7.9 0.712 0.742 4.0 Llama Len 0.834 0.806 -3.5 0.856 0.884 3.1 0.522 0.571 8.7 0.737 0.754 2.8 Llama LogDet 0.817 0.800 -2.1 0.859 0.882 2.6 0.526 0.582 9.6 0.734 0.755 3.4 Llama Mean-Len 0.825 0.798 -3.4 0.852 0.878 2.9 0.509 0.553 7.9 0.729 0.743 2.5 Llama Std-Len 0.820 0.794 -3.2 0.846 0.873 3.1 0.526 0.524 -0.3 0.731 0.730 -0.1 Mistral Perplexity 0.506 0.520 2.7 0.624 0.673 7.4 0.740 0.778 4.9 0.623 0.657 5.0 Mistral LN-Entropy 0.508 0.505 -0.6 0.587 0.615 4.5 0.759 0.825 8.0 0.618 0.648 4.0 Mistral SE 0.872 0.754 -15.7 0.898 0.843 -6.5 0.609 0.538 -13.3 0.793 0.712 -11.8 Mistral Eigenscore 0.873 0.738 -18.4 0.902 0.842 -7.2 0.598 0.567 -5.5 0.791 0.716 -10.4 Mistral eRank 0.515 0.526 2.0 0.855 0.789 -8.4 0.606 0.736 17.8 0.659 0.684 3.8 Mistral Len 0.897 0.735 -22.1 0.918 0.848 -8.3 0.687 0.530 -29.7 0.834 0.704 -20.0 Mistral LogDet 0.895 0.734 -21.9 0.869 0.793 -9.5 0.673 0.561 -19.8 0.812 0.696 -17.1 Mistral Mean-Len 0.879 0.734 -19.7 0.907 0.844 -7.5 0.629 0.548 -14.7 0.805 0.709 -14.0 Mistral Std-Len 0.827 0.683 -20.9 0.873 0.808 -8.0 0.546 0.608 10.1 0.749 0.700 -6.3

### v2

@@ 3.6 Implementation Details
We utilize pretrained model weights from the Hugging Face Transformers (Wolf et al., 2020) without any additional fine-tuning. Following (Farquhar et al., 2024), we generate 10 samples ($n=10$) using temperature $1.0$ for uncertainty estimation. Additionally, we generate one "best answer" sample with temperature $0.1$ to serve as the best-generation estimate for performance evaluation.
@@ 5.1 Limitations of ROUGE for Factual Accuracy Assessment in QA
Model Metric NQ-Open SQuAD Trivia-QA Mean ROUGE LLM $\Delta$% ROUGE LLM $\Delta$% ROUGE LLM $\Delta$% ROUGE LLM $\Delta$% Llama Perplexity 0.709 0.700 -1.2 0.703 0.687 -2.4 0.733 0.789 7.2 0.715 0.725 1.2 LN-Entropy 0.521 0.605 13.9 0.558 0.611 8.7 0.563 0.636 11.5 0.547 0.617 11.4 SE 0.778 0.742 -4.8 0.707 0.705 -0.2 0.769 0.832 7.6 0.751 0.760 0.9 Eigenscore 0.816 0.686 -19.0 0.720 0.638 -12.7 0.752 0.734 -2.5 0.763 0.686 -11.4 eRank 0.825 0.632 -30.6 0.754 0.621 -21.4 0.717 0.660 -8.6 0.765 0.638 -20.2 LogDet 0.511 0.515 0.7 0.521 0.536 2.7 0.604 0.509 -18.6 0.545 0.520 -5.1 Mistral Perplexity 0.852 0.584 -45.9 0.516 0.500 -3.2 0.843 0.627 -34.4 0.737 0.570 -27.8 LN-Entropy 0.718 0.645 -11.3 0.734 0.657 -11.7 0.586 0.596 1.8 0.679 0.633 -7.1 SE 0.836 0.729 -14.7 0.784 0.701 -11.9 0.726 0.707 -2.6 0.782 0.712 -9.7 Eigenscore 0.873 0.669 -30.4 0.803 0.648 -24.0 0.775 0.652 -18.9 0.817 0.656 -24.4 eRank 0.925 0.678 -36.4 0.518 0.511 -1.3 0.851 0.645 -31.9 0.765 0.611 -23.2 LogDet 0.628 0.508 -23.6 0.562 0.518 -8.5 0.843 0.606 -39.2 0.678 0.544 -23.8
@@ 5.3 Impact of Few-Shot Examples on Evaluation Reliability
Few-shot settings consistently yield more reliable evaluations across metrics. For Llama, the discrepancy between ROUGE and LLM-as-Judge narrows significantly with few-shot examples. For instance, eRank’s performance drop (for Llama) reduces from $16.7\%$ in zero-shot to just $4.2\%$ in few-shot settings. This suggests that few-shot examples help standardize response formats with more consistent evaluation.
@@
The impact of few-shot examples varies notably between models. Mistral shows pronounced degradation in zero-shot settings, with performance drops up to $45.9\%$ (Perplexity), while Llama maintains more consistent performance, with some metrics showing minimal degradation. This variation suggests that the architecture and pre-training may influence the effectiveness of few-shot calibration.
@@ G.1 Length-Based Penalties
ROUGE systematically penalizes factually correct but verbose answers. In this example, despite providing accurate information with helpful context, the response receives a low score purely due to length mismatch. As shown in Figure 5, this bias affects longer responses regardless of their factual accuracy, with responses exceeding $100$ tokens consistently scoring below our $0.3$ threshold. Notably, this is the most frequent type of error ROUGE makes.
@@ H.2 Metric Evaluation: AUROC
Model Metric NQ-Open SQuAD Trivia-QA Mean ROUGE LLM $\Delta$% ROUGE LLM $\Delta$% ROUGE LLM $\Delta$% ROUGE LLM $\Delta$% Llama Perplexity 0.709 0.700 -1.2 0.703 0.687 -2.4 0.733 0.789 7.2 0.715 0.725 1.2 Llama LN-Entropy 0.521 0.605 13.9 0.558 0.611 8.7 0.563 0.636 11.5 0.547 0.617 11.4 Llama SE 0.778 0.742 -4.8 0.707 0.705 -0.2 0.769 0.832 7.6 0.751 0.760 0.9 Llama Eigenscore 0.816 0.686 -19.0 0.720 0.638 -12.7 0.752 0.734 -2.5 0.763 0.686 -11.4 Llama eRank 0.825 0.632 -30.6 0.754 0.621 -21.4 0.717 0.660 -8.6 0.765 0.638 -20.2 Llama Len 0.834 0.616 -35.3 0.777 0.622 -24.9 0.760 0.691 -10.0 0.790 0.643 -23.4 Llama LogDet 0.511 0.515 0.7 0.521 0.536 2.7 0.604 0.509 -18.6 0.545 0.520 -5.1 Llama Mean-Len 0.825 0.654 -26.1 0.743 0.643 -15.7 0.771 0.743 -3.8 0.780 0.680 -15.2 Llama Std-Len 0.711 0.644 -10.5 0.664 0.627 -6.0 0.759 0.754 -0.7 0.711 0.675 -5.7 Mistral Perplexity 0.852 0.584 -45.9 0.516 0.500 -3.2 0.843 0.627 -34.4 0.737 0.570 -27.8 Mistral LN-Entropy 0.718 0.645 -11.3 0.734 0.657 -11.7 0.586 0.596 1.8 0.679 0.633 -7.1 Mistral SE 0.836 0.729 -14.7 0.784 0.701 -11.9 0.726 0.707 -2.6 0.782 0.712 -9.7 Mistral Eigenscore 0.873 0.669 -30.4 0.803 0.648 -24.0 0.775 0.652 -18.9 0.817 0.656 -24.4 Mistral eRank 0.925 0.678 -36.4 0.518 0.511 -1.3 0.851 0.645 -31.9 0.765 0.611 -23.2 Mistral Len 0.934 0.634 -47.2 0.860 0.624 -37.8 0.929 0.673 -37.9 0.908 0.644 -41.0 Mistral LogDet 0.628 0.508 -23.6 0.562 0.518 -8.5 0.843 0.606 -39.2 0.678 0.544 -23.8 Mistral Mean-Len 0.890 0.643 -38.4 0.828 0.626 -32.2 0.875 0.667 -31.3 0.864 0.645 -34.0 Mistral Std-Len 0.516 0.512 -0.7 0.540 0.505 -6.9 0.613 0.572 -7.2 0.556 0.530 -4.9
@@
Model Metric NQ-Open SQuAD Trivia-QA Mean ROUGE LLM $\Delta$% ROUGE LLM $\Delta$% ROUGE LLM $\Delta$% ROUGE LLM $\Delta$% Llama Perplexity 0.814 0.767 -6.1 0.736 0.758 2.9 0.800 0.826 3.1 0.783 0.784 -0.0 Llama LN-Entropy 0.753 0.732 -2.9 0.663 0.717 7.5 0.799 0.829 3.6 0.738 0.759 2.7 Llama SE 0.738 0.730 -1.1 0.688 0.741 7.1 0.800 0.849 5.7 0.742 0.773 3.9 Llama Eigenscore 0.813 0.744 -9.3 0.725 0.733 1.2 0.745 0.762 2.3 0.761 0.746 -1.9 Llama eRank 0.794 0.714 -11.2 0.708 0.681 -4.0 0.620 0.638 2.8 0.707 0.678 -4.1 Llama Len 0.761 0.686 -10.9 0.694 0.687 -1.0 0.620 0.640 3.1 0.692 0.671 -2.9 Llama LogDet 0.729 0.690 -5.6 0.659 0.636 -3.7 0.590 0.618 4.5 0.659 0.648 -1.6 Llama Mean-Len 0.799 0.730 -9.4 0.713 0.716 0.4 0.681 0.716 4.8 0.731 0.721 -1.4 Llama Std-Len 0.777 0.727 -7.0 0.705 0.721 2.2 0.783 0.806 2.9 0.755 0.751 -0.6 Mistral Perplexity 0.804 0.632 -27.1 0.782 0.636 -23.0 0.744 0.637 -16.7 0.777 0.635 -22.3 Mistral LN-Entropy 0.727 0.619 -17.4 0.785 0.667 -17.7 0.750 0.692 -8.3 0.754 0.659 -14.5 Mistral SE 0.772 0.734 -5.3 0.737 0.698 -5.6 0.741 0.765 3.1 0.750 0.732 -2.6 Mistral Eigenscore 0.789 0.686 -15.0 0.775 0.691 -12.2 0.717 0.706 -1.5 0.760 0.694 -9.6 Mistral eRank 0.874 0.698 -25.1 0.829 0.690 -20.1 0.786 0.703 -11.8 0.830 0.697 -19.0 Mistral Len 0.879 0.664 -32.2 0.857 0.685 -25.1 0.858 0.729 -17.7 0.865 0.693 -25.0 Mistral LogDet 0.737 0.663 -11.2 0.687 0.631 -8.9 0.612 0.630 2.9 0.679 0.641 -5.7 Mistral Mean-Len 0.834 0.683 -22.1 0.822 0.705 -16.5 0.806 0.750 -7.4 0.821 0.713 -15.3 Mistral Std-Len 0.609 0.577 -5.6 0.629 0.589 -6.8 0.663 0.665 0.3 0.634 0.610 -4.0
@@ H.3 Metric Evaluation: PR-AUC
Model Metric NQ-Open SQuAD Trivia-QA Mean ROUGE LLM $\Delta$% ROUGE LLM $\Delta$% ROUGE LLM $\Delta$% ROUGE LLM $\Delta$% Llama Perplexity 0.833 0.680 -22.4 0.863 0.823 -4.8 0.594 0.514 -15.6 0.763 0.672 -14.3 Llama LN-Entropy 0.717 0.611 -17.4 0.793 0.773 -2.6 0.570 0.652 12.5 0.693 0.679 -2.5 Llama SE 0.845 0.695 -21.5 0.864 0.829 -4.1 0.575 0.533 -7.9 0.761 0.686 -11.2 Llama Eigenscore 0.850 0.670 -26.8 0.866 0.809 -7.1 0.565 0.574 1.6 0.760 0.684 -10.8 Llama eRank 0.782 0.607 -28.9 0.820 0.783 -4.6 0.674 0.760 11.3 0.759 0.717 -7.4 Llama Len 0.865 0.681 -27.2 0.885 0.820 -8.0 0.605 0.548 -10.4 0.785 0.683 -15.2 Llama LogDet 0.852 0.659 -29.2 0.873 0.810 -7.8 0.602 0.562 -7.1 0.776 0.677 -14.7 Llama Mean-Len 0.851 0.658 -29.3 0.870 0.808 -7.7 0.573 0.568 -0.9 0.765 0.678 -12.6 Llama Std-Len 0.825 0.647 -27.6 0.846 0.802 -5.5 0.562 0.570 1.4 0.744 0.673 -10.6 Mistral Perplexity 0.664 0.536 -23.8 0.951 0.754 -26.0 0.690 0.752 8.3 0.768 0.681 -13.8 Mistral LN-Entropy 0.882 0.664 -32.8 0.920 0.790 -16.4 0.625 0.633 1.3 0.809 0.696 -16.0 Mistral SE 0.956 0.725 -31.8 0.964 0.819 -17.7 0.808 0.510 -58.3 0.909 0.685 -35.9 Mistral Eigenscore 0.957 0.698 -37.1 0.965 0.804 -20.0 0.818 0.544 -50.3 0.913 0.682 -35.8 Mistral eRank 0.658 0.506 -30.0 0.955 0.755 -26.4 0.534 0.704 24.2 0.716 0.655 -10.7 Mistral Len 0.964 0.682 -41.4 0.973 0.803 -21.0 0.849 0.536 -58.4 0.929 0.674 -40.3 Mistral LogDet 0.964 0.699 -37.9 0.950 0.753 -26.1 0.847 0.550 -54.1 0.920 0.667 -39.4 Mistral Mean-Len 0.958 0.671 -42.8 0.966 0.786 -22.9 0.833 0.548 -52.1 0.919 0.668 -39.3 Mistral Std-Len 0.891 0.583 -52.8 0.889 0.724 -22.7 0.755 0.605 -24.8 0.845 0.637 -33.4
@@
Model Metric NQ-Open SQuAD Trivia-QA Mean ROUGE LLM $\Delta$% ROUGE LLM $\Delta$% ROUGE LLM $\Delta$% ROUGE LLM $\Delta$% Llama Perplexity 0.844 0.824 -2.4 0.861 0.891 3.4 0.551 0.502 -9.8 0.752 0.739 -2.9 Llama LN-Entropy 0.810 0.796 -1.8 0.828 0.874 5.3 0.525 0.522 -0.5 0.721 0.731 1.0 Llama SE 0.814 0.802 -1.5 0.842 0.879 4.3 0.536 0.506 -6.1 0.731 0.729 -1.1 Llama Eigenscore 0.829 0.802 -3.4 0.852 0.876 2.7 0.511 0.542 5.7 0.731 0.740 1.7 Llama eRank 0.746 0.726 -2.8 0.711 0.762 6.8 0.679 0.737 7.9 0.712 0.742 4.0 Llama Len 0.834 0.806 -3.5 0.856 0.884 3.1 0.522 0.571 8.7 0.737 0.754 2.8 Llama LogDet 0.817 0.800 -2.1 0.859 0.882 2.6 0.526 0.582 9.6 0.734 0.755 3.4 Llama Mean-Len 0.825 0.798 -3.4 0.852 0.878 2.9 0.509 0.553 7.9 0.729 0.743 2.5 Llama Std-Len 0.820 0.794 -3.2 0.846 0.873 3.1 0.526 0.524 -0.3 0.731 0.730 -0.1 Mistral Perplexity 0.506 0.520 2.7 0.624 0.673 7.4 0.740 0.778 4.9 0.623 0.657 5.0 Mistral LN-Entropy 0.508 0.505 -0.6 0.587 0.615 4.5 0.759 0.825 8.0 0.618 0.648 4.0 Mistral SE 0.872 0.754 -15.7 0.898 0.843 -6.5 0.609 0.538 -13.3 0.793 0.712 -11.8 Mistral Eigenscore 0.873 0.738 -18.4 0.902 0.842 -7.2 0.598 0.567 -5.5 0.791 0.716 -10.4 Mistral eRank 0.515 0.526 2.0 0.855 0.789 -8.4 0.606 0.736 17.8 0.659 0.684 3.8 Mistral Len 0.897 0.735 -22.1 0.918 0.848 -8.3 0.687 0.530 -29.7 0.834 0.704 -20.0 Mistral LogDet 0.895 0.734 -21.9 0.869 0.793 -9.5 0.673 0.561 -19.8 0.812 0.696 -17.1 Mistral Mean-Len 0.879 0.734 -19.7 0.907 0.844 -7.5 0.629 0.548 -14.7 0.805 0.709 -14.0 Mistral Std-Len 0.827 0.683 -20.9 0.873 0.808 -8.0 0.546 0.608 10.1 0.749 0.700 -6.3